Which equation gives you a solution of (2,0) ?
A. 3x - y = 6; y + 2 = x
B. 3x + y = 6; y + 2 = x
C. 3x - y = 6; y + 2 = -x
D. 3x + y = 6; y + 2 = -x

Answer by xinxin(76)

The easiest way for solving this kind of problem is to plug given answers into each choice, and see if you can get correct equations:
A: 3*2-0=6, 0+2=2
B: 3*2+0=6, 0+2=2
C&D are incorrect. So you should choose A and B.